OpenStudy (anonymous):

not sure where to start for this proof. Determine whether the following sets form subspaces of R^2 {(x1,x2)^T | x1+x2=0}

OpenStudy (jtvatsim):

Do you know the conditions required to have a subspace?

OpenStudy (anonymous):

it has to be nonempty alpha x is in the subset for any alpha x+y is in the subset that is what my book says. but not sure how to apply that to this

OpenStudy (jtvatsim):

OK, there are three properties as your book suggests. Another equivalent version says: 1) the zero vector must be in the subset (this ensures nonemptiness, as well as other useful things) 2) constant multiples must be in the subset (alpha x, where alpha is a scalar and x is a vector) 3) vector sums must be in the subset (x + y, where x and y are both vectors).

OpenStudy (jtvatsim):

The trick, as you mention, is figuring out how to apply it. :)

OpenStudy (jtvatsim):

We are dealing with the subset {(x1,x2)^T | x1 + x2 = 0} correct? Are you clear on what all the notation means here?

OpenStudy (anonymous):

i think one of the problems is that calculus 1 was the only pre req for linear but trying to read the book i think its assumed you have had discrete or some other math proof based class. at least thats what i get from talling to other students

OpenStudy (anonymous):

talking*

OpenStudy (jtvatsim):

Yes, you are probably correct... I imagine it is quite a bit more challenging in that case... :) It is easy enough once you get the basics. I'll try to walk you through this one if you think that will be most helpful.

OpenStudy (anonymous):

anything would be. im lost and tried googling stuff but its just more proofs and symbols that i then have to try and google what those symboles mean

OpenStudy (jtvatsim):

OK. Let's start with step one. Identify what all the important concepts, symbols, and definitions are.

OpenStudy (anonymous):

ok. i know that x1 and x2 are vectors in R^2

OpenStudy (anonymous):

and im assuming R^2 is the x y plane

OpenStudy (anonymous):

and there sum is 0

OpenStudy (jtvatsim):

Good, so far. There is a minor issue however.

OpenStudy (freckles):

well I think (x1,x2)^T is a vector in R^2 not x1 and x2

OpenStudy (anonymous):

thats about as far as i can go

OpenStudy (anonymous):

what issue?

OpenStudy (jtvatsim):

That's a good start. As @freckles mentioned, we need to be careful to distinguish that x1 and x2 are not vectors separately.

OpenStudy (jtvatsim):

The whole expression is (x1,x2) which is a vector. x1 is the x-component, and x2 is the y-component. These individual components sum to 0.

OpenStudy (anonymous):

so its a single vector like x = (x1 x2)

OpenStudy (jtvatsim):

Correct

OpenStudy (jtvatsim):

It is useful to think of an actual example just so we have something concrete in mind. So some examples of vectors that satisfy this are (1,-1) since 1 + (-1) = 0, (-2, 2) since (-2) + 2 = 0. and so on. Make sense so far?

OpenStudy (anonymous):

yes. i guess "abstractly" x1 = negative x2 so that they sum to 0

OpenStudy (jtvatsim):

Exactly! In fact, that is what the expression x1 + x2 = 0 implies. Solving for x1 gives as you say, x1 = -x2.

OpenStudy (jtvatsim):

Alright, so we still need to decide whether this set of vectors is in fact a subspace.

OpenStudy (jtvatsim):

Let's start with the first condition: 1) Does the zero vector live in this set?

OpenStudy (anonymous):

i would say yes because 0 + (-0) is still 0. is that the right way to think of it?

OpenStudy (jtvatsim):

You are 100% correct!

OpenStudy (jtvatsim):

(0,0) satisfies the condition that x1 + x2 = 0, since 0 + 0 = 0. Good, our set passes condition 1.

OpenStudy (jtvatsim):

Now, to check condition 2: If c is some multiple and v is in the set, is cv in the set?

OpenStudy (jtvatsim):

Sorry, I guess I should say If alpha is some multiple and x is in the set, is alpha x in the set?

OpenStudy (jtvatsim):

It's the same thing either way.

OpenStudy (jtvatsim):

So, this is a little weird, but here goes.

OpenStudy (anonymous):

so alphax1 + alphax2 = 0 so alpha is any number alpha = 100 x1 = 5 so x2 would = -5 so someting like 100(5) + 100(-5) = 0 because 500 + -500 = 0

OpenStudy (jtvatsim):

Great, that's the spirit of the proof! Now we just make the argument as general as possible.

OpenStudy (anonymous):

thats where i run into trouble.

OpenStudy (anonymous):

becasue i asked him if i can give an example on a test that is coming up and he said no because one example might be right but there is posibly infinite examples and one could be false and some other stuff i didnt quite catch

OpenStudy (jtvatsim):

Here is how to phrase the general proof. Look it over and if you have any questions, we will spend some time unpacking it. "Suppose we have x = (x1,x2) in the set. So, x1 + x2 = 0. Now, look at alpha x. alpha x = (alpha x1, alpha x2). Notice that alpha x1 + alpha x2 = alpha (x1 + x2). But we know that x1 + x2 = 0! So, alpha x1 + alpha x2 = alpha (x1 + x2) = 0 as we wanted."

OpenStudy (jtvatsim):

Notice that we need to start with x. We do not know if alpha x1 + alpha x2 actually equals 0. This is what we are trying to show... Instead we start with x and show that alpha x works too. With a couple of clever algebraic tricks we pull it off.

OpenStudy (anonymous):

ok i think i get it, one question though, might be dumb but where you (and my book) say constant multiples must be in the subset. lets say the subset is 1 2 3 4 5 does that mean alpha can only be 1 2 3 4 5 and not 6 or anything higher or am i completely wrong on that

OpenStudy (anonymous):

but what you did is only for the second condition, right?

OpenStudy (jtvatsim):

This is only the second condition, yes. And your question is a good one. Let me think how to answer it.

OpenStudy (jtvatsim):

alpha can be as large or as small as it wants. It must work for any alpha. So, in your example, you give a very small subset. {1, 2, 3, 4, 5}. This is actually NOT a subspace since alpha = 10 breaks it. In fact 10*1 = 10 which is not in the set.

OpenStudy (jtvatsim):

Does that answer your question?

OpenStudy (anonymous):

ok, yes it does

OpenStudy (jtvatsim):

Great. One last note on condition 2. We have to start with x1 + x2 = 0 NOT alpha x1 + alpha x2 = 0 since the set only talks about x1 + x2 = 0 in the definition.

OpenStudy (jtvatsim):

That was a minor difference in your proof versus my general proof.

OpenStudy (jtvatsim):

Just wanted to be sure you picked up on that. :)

OpenStudy (jtvatsim):

So where are we so far? We have shown that condition 1 and 2 both work. All that remains is condition 3. If it works, then we have a subspace. If it fails, we do not have a subspace.

OpenStudy (anonymous):

i scrolled up and read what you said the three conditinos are and number 3 says someting about 2 vectors but dont be only have 1 vector x ?

OpenStudy (jtvatsim):

Good point. We are allowed to assume we have a second vector y for condition 3's proof.

OpenStudy (jtvatsim):

After all, with condition 2 we just showed that this set probably has an infinite number of vectors (every multiple works).

OpenStudy (anonymous):

so would the second vector consist of 0's ?

OpenStudy (jtvatsim):

Well, we don't exactly know. It is safer to say that the second vector y = (y1,y2) will just act like the other vectors in this set. That is, y1 + y2 = 0.

OpenStudy (anonymous):

because anything added to 0 is still whatever that number was. so 5 + 0 = 5?

OpenStudy (anonymous):

is that right or wrong way to think of it?

OpenStudy (anonymous):

ok so then again y1 = -y2

OpenStudy (jtvatsim):

I see your point. However, your example is dealing with numbers rather than vectors.

OpenStudy (jtvatsim):

Yes, that is the safer assumption.

OpenStudy (jtvatsim):

But hold your thought, you might need just a slight change in thinking to see what we are getting at.

OpenStudy (jtvatsim):

So, how will we know if x + y is in the set?

OpenStudy (jtvatsim):

We need to show that its components add to 0.

OpenStudy (anonymous):

if the vectors of x = -y ?

OpenStudy (jtvatsim):

Yes, I think you could do it that way.

OpenStudy (anonymous):

is that not the right way? or was there something else?

OpenStudy (jtvatsim):

Ah... wait. We aren't saying that x + y = 0. We want to say that (first components) + (second components) = 0.

OpenStudy (jtvatsim):

Here's what I mean.

OpenStudy (jtvatsim):

First, we need to know what x + y looks like. x + y = (x1,x2) + (y1,y2) = (x1+y1,x2+y2)

OpenStudy (jtvatsim):

Then, we need to find a way to show that x1+y1 + x2+y2 = 0. That will be the trick.

OpenStudy (jtvatsim):

Does that make sense? Any questions on that part?

OpenStudy (anonymous):

i think so. wouldt you just move x2 and y2 over? so x1 + x2 = -x2 - y2 ??

OpenStudy (anonymous):

actually i think thats wrong

OpenStudy (jtvatsim):

You could do that. But that won't be a proof. We need to start with the most basic assumption and build up to what we want. So here we go (*long inhale*)... :)

OpenStudy (jtvatsim):

"Suppose we have two vectors x and y in the set."

OpenStudy (jtvatsim):

"Then, x = (x1,x2) so x1 + x2 = 0 and y = (y1,y2) so y1 + y2 = 0. This is true since they are in the set."

OpenStudy (jtvatsim):

"Now, I'd like to think about x + y. x + y looks like this: x + y = (x1,x2) + (y1,y2) = (x1+y1,x2+y2)."

OpenStudy (jtvatsim):

"Now, do these components sum to 0? I notice that (x1 + y1) + (x2 + y2) = x1 + y1 + x2 + y2 = (x1 + x2) + (y1 + y2)"

OpenStudy (jtvatsim):

"But, wait, I just said before that x1 + x2 = 0 and y1 + y2 = 0. So, the components do sum to 0 after all! In fact, (x1 + y1) + (x2 + y2) = (x1 + x2) + (y1 + y2) = 0 + 0 = 0 as I wanted."

OpenStudy (jtvatsim):

Do you follow? There are little nuances here and there, but this is really the same format you follow for any "subspace proof". :)

OpenStudy (jtvatsim):

I'll make a summary of the highlights so you can bring all these scattered ideas together. :)

OpenStudy (anonymous):

ok. im still reading everything you just typed and comprehending, might take a minute

OpenStudy (jtvatsim):

No worries, take your time, these things don't happen in a flash. It took me a lot of practice to get comfortable. :)

OpenStudy (jtvatsim):

This is what a finished proof might look like (you might find it helpful to see the big picture all at once): {(x1,x2)^T | x1 + x2 = 0} is a subspace of R^2. Proof: A set is a subspace if it is 1) nonempty, 2) cx is in the set for any scalar c, and 3) if x + y is in the set. 1) (0,0) is in the set since 0 + 0 = 0. So, the set is nonempty. 2) Suppose x is in the set and c is some scalar. Then x = (x1,x2) and x1 + x2 = 0. Think about cx. cx = (cx1, cx2). Now cx1 + cx2 = c(x1 + x2) = c(0) = 0. So cx is in the set. 3) Suppose x and y are in the set. Then x = (x1,x2) and x1 + x2 = 0, also y = (y1,y2) and y1 + y2 = 0. Think about x + y. x + y = (x1,x2) + (y1,y2) = (x1+y1,x2+y2). Now (x1 + y1) + (x2 + y2) = (x1 + x2) + (y1 + y2) = 0 + 0 = 0. So x + y is in the set. Since all three conditions are met, the set is a subspace.

OpenStudy (anonymous):

that was both kind of morerinvolved than i was thinking but not that bad once you explained it. im still reading. but it is making sense

OpenStudy (jtvatsim):

Yes, proofs are quite picky, but the format is almost always the same.

OpenStudy (jtvatsim):

The most frustrating thing is not being able to just start with the goal. We cannot say: "Suppose alpha x1 + alpha x2 = 0..." and go from there. We must say: "Suppose x is in the set. Then x1 + x2 = 0..." and somehow manipulate this to get to "alpha x1 + alpha x2 = 0". It's a subtle but crucial difference in direction.

OpenStudy (anonymous):

ok. i have some similar problens to work on. like x1*x2 = 0 and a few others but im going to try and apply what you explained to them. i might hit you up if i get stuck but im going to try them first

OpenStudy (jtvatsim):

Sounds good. Remember that not all sets will be subspaces. If you an find a way to break any of the three rules you are allowed to say "No way, this is not a subspace."

OpenStudy (jtvatsim):

Also a "Pro Tip," subspaces are linear, so if you have multiplication of variables (like x1*x2) you will probably not have a subspace.

OpenStudy (anonymous):

so is there a different way to work it?

OpenStudy (jtvatsim):

If you don't have a subspace, just write down the example that breaks it. Are you working on the x1*x2 = 0 one?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, I'm looking at it now.

OpenStudy (jtvatsim):

Well, let's just go through the list. Does (0,0) work?

OpenStudy (anonymous):

so is x = (1,0)

OpenStudy (anonymous):

or x = (0,1)

OpenStudy (jtvatsim):

Yep, all those are in the set so far.

OpenStudy (anonymous):

well yes 0 *0 =0